### 14 May 2026 v1

# Boundary-supported radial layering in Hoag-like ring galaxies

Abstract
Hoag-like galaxies exhibit an old compact central component, a depleted gap, and a detached outer ring. We identify a boundary-supported radial-layering mechanism in a shell-deformed Kepler control model. A compact inner boundary supplies the core state, while a localized shell, interpreted as the reduced imprint of externally supplied material settled near a finite circularization radius, needs to create only an internal maximum and a subsequent outer minimum. These act as the gap barrier and ring-supporting well. The onset of this structure is organized by a saddle-node threshold of the critical-point equation. In a $10^{4}$-point Monte Carlo scan, shell-localized boundary-supported candidates occupy finite parameter volume, and none of the localized candidates contains an ordered interior minimum–maximum–minimum subsequence. The same branch gives a scale-free gap-to-ring interval consistent with Hoag’s Object, UGC 4599, and PGC 1000714, but not with the environmentally processed comparison object JO171.
keywords:
galaxies: peculiar – galaxies: structure – galaxies: evolution – galaxies: kinematics and dynamics
pubyear: 2026pagerange: Boundary-supported radial layering in Hoag-like ring galaxiesReferences
Problem and positioning.
Hoag-like ring galaxies are distinguished not simply by the presence of a ring, but by a clean radial separation: a compact central body, an underdense gap, and a detached outer annulus. This structure is difficult to reduce to the standard ring-formation channels. Nearly head-on collisions generate expanding density waves (Lynds & Toomre, 1976; Appleton & Struck-Marcell, 1996), while bar or oval perturbations organize gas near resonances in disk galaxies (Buta & Combes, 1996). The cleanest Hoag-type systems, however, often show spheroidal or nearly round cores, weak or absent bar signatures, and evidence for externally supplied gas or extended H i reservoirs (Schweizer et al., 1987; Finkelman et al., 2011; Finkelman & Brosch, 2011; Pakdil et al., 2017; Moretti et al., 2018).
These observations motivate a more specific structural question. If an old compact core is already present and later accreted gas dissipates and settles near a finite radius, what is the minimal radial geometry needed to maintain a depleted interval between the core and the outer ring? A related mechanical problem was studied by Bannikova, who considered a central mass plus a massive ring or torus and showed that the competition of their gravitational fields can create an OSCO–LC region where stable circular motion is absent (Bannikova, 2018). Our mechanism is complementary but different. We do not require the ring to be a fully self-gravitating torus. Instead we ask whether the reduced radial imprint of a settled shell can create a barrier–well structure outside a compact inner boundary.
The answer is yes, but the topology is not the usual interior double well. In a regular effective potential, two detached zones would be represented by
${\rm minimum}\;-\;{\rm maximum}\;-\;{\rm minimum},$ (1)
with both the core and the ring generated by internal critical points. For a compact Kepler-like background the inner state is instead supplied by the boundary. The required sequence becomes
${\rm boundary}\;-\;{\rm maximum}\;-\;{\rm minimum}.$ (2)
This distinction is the central result of the Letter: Hoag-like radial layering can be organized by a boundary-supported shell-localized bifurcation rather than by a fully interior double well.
Shell-deformed Kepler control model.
Let
$x=\frac{r}{R_{\rm s}}$ (3)
be a dimensionless radius. The control potential is
$\Phi(x)=-\frac{\mu}{x}+\epsilon\exp\!\left[-\frac{(x-x_{0})^{2}}{2\Delta^{2}}\right],\qquad\mu>0.$ (4)
The Kepler term represents the compact-core control limit. The Gaussian term is an effective reduced deformation, not a Poisson-self-consistent gas potential; it represents the radial imprint of externally supplied material that has dissipated, partially circularized, and accumulated near a finite radius. Thus $x_{0}$, $\Delta$, and $\epsilon$ describe the settling radius, radial width, and control strength of the shell.
We diagnose radial barriers and wells using
$W(x;\ell)=e^{2\Phi(x)}\left(1+\frac{\ell^{2}}{x^{2}}\right),$ (5)
where $\ell$ labels the angular-momentum scale of material participating in the shell-supported structure. This form can be motivated by an auxiliary radial geodesic construction. For the effective line element
$ds^{2}=-e^{2\Phi(x)}dt^{2}+e^{-2\Phi(x)}dx^{2}+x^{2}d\varphi^{2},$ (6)
timelike curves have conserved labels $E=e^{2\Phi}\dot{t}$ and $\ell=x^{2}\dot{\varphi}$, and the normalization condition gives
$\dot{x}^{2}+e^{2\Phi(x)}\left(1+\frac{\ell^{2}}{x^{2}}\right)=E^{2}.$ (7)
Equation (5) is therefore a reduced square-energy function. It is used only as an effective radial control diagnostic; in the weak-field limit
$W=1+2\Phi+\frac{\ell^{2}}{x^{2}}+O\!\left(\Phi^{2},\frac{\Phi\ell^{2}}{x^{2}}\right),$ (8)
so its extrema track the leading Newtonian barrier–well ordering.
Critical radii satisfy $W^{\prime}(x_{\ast})=0$, equivalently
$\frac{d\Phi}{dx}=\frac{\ell^{2}}{x(x^{2}+\ell^{2})}.$ (9)
A local maximum of $W$ is interpreted as a radial barrier and a local minimum as a supporting well. The Kepler boundary gives
$W(x)\rightarrow 0\quad(x\rightarrow 0),\qquad W(x)\rightarrow 1\quad(x\rightarrow\infty).$ (10)
Thus the central core state is not an interior critical point of $W(x)$; it is represented by the compact inner boundary. The shell needs to create only the gap barrier and the outer ring well.
Accretion-settling interpretation.
The shell deformation can be embedded in a minimal dissipative accretion picture. Gas supplied from large radii is labeled by an angular- momentum scale $\ell_{i}$ and evolves, in the weak-field limit, in the instantaneous effective potential
$U_{\rm eff}(x;\ell_{i},\tau)=\Phi(x,\tau)+\frac{\ell_{i}^{2}}{2x^{2}}.$ (11)
A reduced radial equation is
$\frac{d^{2}x_{i}}{d\tau^{2}}+\Gamma\frac{dx_{i}}{d\tau}=-\partial_{x}U_{\rm eff}(x_{i};\ell_{i},\tau),$ (12)
where $\Gamma>0$ represents radial-energy loss through gas collisions, radiative cooling, shocks, and turbulent stresses. The radial energy
$E_{r,i}=\frac{1}{2}\left(\frac{dx_{i}}{d\tau}\right)^{2}+U_{\rm eff}(x_{i};\ell_{i},\tau)$ (13)
obeys
$\frac{dE_{r,i}}{d\tau}=-\Gamma\left(\frac{dx_{i}}{d\tau}\right)^{2}+\partial_{\tau}U_{\rm eff}.$ (14)
Thus, when the shell grows slowly compared with the damping time, gas in the outer basin loses radial energy and approaches the ring-supporting minimum, while the intervening maximum remains a low-residence barrier. In this sense the static max–min structure identified below is the instantaneous barrier–attractor landscape for later dissipative settling.
Boundary-supported saddle-node threshold.
Writing Eq. (9) as $F(x)=0$ gives
$F(x)=A(x)-\epsilon G(x),$ (15)
where
$A(x)=\frac{\mu}{x^{2}}-\frac{\ell^{2}}{x(x^{2}+\ell^{2})}$ (16)
and
$G(x)=\frac{x-x_{0}}{\Delta^{2}}\exp\!\left[-\frac{(x-x_{0})^{2}}{2\Delta^{2}}\right].$ (17)
The birth of a shell-induced maximum–minimum pair is a saddle-node threshold of this one-dimensional critical-point equation (Strogatz, 2015; Arnold, 1986). At threshold,
$F(x_{\rm c})=0,\qquad F^{\prime}(x_{\rm c})=0,$ (18)
or, equivalently,
$A^{\prime}(x_{\rm c})G(x_{\rm c})-A(x_{\rm c})G^{\prime}(x_{\rm c})=0,\qquad\epsilon_{\rm c}=\frac{A(x_{\rm c})}{G(x_{\rm c})}.$ (19)
The Monte Carlo branch below is therefore organized by an analytic threshold surface rather than by an accidental numerical root pattern.
After the pair forms, let
$x_{\rm g}<x_{\rm r}$ (20)
denote the maximum and subsequent minimum:
$\begin{gathered}W^{\prime}(x_{\rm g})=0,\quad W^{\prime\prime}(x_{\rm g})<0,\\ W^{\prime}(x_{\rm r})=0,\quad W^{\prime\prime}(x_{\rm r})>0.\end{gathered}$ (21)
The maximum $x_{\rm g}$ is the gap barrier and the minimum $x_{\rm r}$ is the ring-supporting well. On a finite grid we define $W_{\rm in}=W(x_{\min})$, $W_{\rm out}=W(x_{\max})$, $W_{\rm g}=W(x_{\rm g})$, and $W_{\rm r}=W(x_{\rm r})$. A candidate is boundary-supported if
$W_{\rm in}<W_{\rm r}<W_{\rm g},\qquad W_{\rm r}<W_{\rm out}.$ (22)
To ensure that the pair is genuinely driven by the shell, we impose
$0.8\leq\frac{x_{\rm g}}{x_{0}}\leq 1.2,\qquad 1.0\leq\frac{x_{\rm r}}{x_{0}}\leq 2.5.$ (23)
The singular Kepler form should be understood as the compact-core limit. A regularized core,
$\Phi_{\rm c,reg}(x)=-\frac{\mu}{(x^{2}+x_{\rm c}^{2})^{1/2}},$ (24)
differs from $\mu/x$ by $O(x_{\rm c}^{2}/x^{3})$ outside the unresolved central region. Since accepted shells satisfy $x_{\rm g}\sim x_{0}$ with $x_{0}=O(1)$, the shell-induced roots are perturbatively unchanged for $x_{\rm c}\ll x_{\rm g}$. The boundary condition is then imposed at an inner numerical radius $x_{\rm in}\sim x_{\rm c}$ by requiring $W(x_{\rm in})<W_{\rm r}$. This formulation makes clear that the mechanism depends on a compact inner boundary, not on treating the mathematical singularity as a literal galactic core.
Figure 1: Boundary-supported radial-layering mechanism. The compact inner boundary supplies the core state. A localized shell centered at $x_{0}$ creates an internal maximum $x_{\rm g}$ and an outer minimum $x_{\rm r}$, identified with the gap barrier and ring-supporting well.
Alt text: Schematic radial profile showing a compact inner boundary, a shell-centered maximum labelled as the gap barrier, and an outer minimum labelled as the ring-supporting well.
Monte Carlo evidence.
We scanned the five-dimensional parameter space
$(\mu,\ell,x_{0},\Delta,\epsilon)$ (25)
using $10^{4}$ random Monte Carlo realizations. The ranges were
$\begin{gathered}\mu\in[0.1,3],\quad\ell\in[0.3,3],\quad x_{0}\in[2,8],\\ \Delta\in[0.1,1],\quad\epsilon\in[0,2],\end{gathered}$ (26)
with log-uniform sampling for $\mu$, $\ell$, and $\Delta$, uniform sampling for $x_{0}$ and $\epsilon$, and $\Delta/x_{0}\leq 0.5$. For each realization all internal critical points of $W$ were located and classified.
Interior minimum–maximum–minimum subsequences can occur in the full sample, usually as part of profiles with additional inner structure. However, they are disjoint from the shell-localized Hoag-like branch: none of the localized candidates contains an ordered interior minimum–maximum–minimum subsequence. By contrast, internal maximum–minimum pairs occurred with fraction
$f_{\rm max-min}=0.5551,$ (27)
and every such realization satisfied the boundary inequalities in Eq. (22) within the sampled domain. After applying the localization condition in Eq. (23), we obtained
$N_{\rm loc}=5494,\qquad f_{\rm loc}=0.5494.$ (28)
Thus the dominant Hoag-like branch is not an ordinary interior double well hidden in a small corner of parameter space. It is a shell-localized maximum–minimum pair supported by the compact inner boundary.
Figure 2 shows the corresponding localization diagnostic. Accepted candidates cluster near $x_{\rm g}/x_{0}\simeq 1$, demonstrating that the gap barrier tracks the settled shell. The ring well lies outside the shell, typically with $x_{\rm r}/x_{0}>1$. This behavior is the key difference from a generic curve-fitting potential: the model identifies where the gap should be relative to the shell.
Figure 2: Monte Carlo evidence for shell localization. Boundary-supported candidates cluster near $x_{\rm g}/x_{0}\simeq 1$, while the ring-supporting minimum forms outside the shell. Sparse nonlocal roots are removed by Eq. (23).
Alt text: Scatter plot of accepted Monte Carlo candidates in shell-localization coordinates, concentrated near gap-to-shell ratio one with ring-to-shell ratio greater than one.
Observable scaling.
For each localized candidate, an observed ring radius fixes
$R_{\rm s}=\frac{R_{\rm ring}^{\rm obs}}{x_{\rm r}}.$ (29)
The model then gives
$\frac{R_{\rm gap}}{R_{\rm ring}}=\frac{x_{\rm g}}{x_{\rm r}},$ (30)
which is independent of the absolute distance scale. We define $R_{\rm ring}$ as the outer-ring peak radius, or the representative mid-ring radius when a peak is not tabulated, and $R_{\rm gap}$ as the surface-brightness minimum between core and ring, or the inner onset of the detached ring when that minimum is not separately reported. Because these literature radii are heterogeneous, the comparison below is a morphology-scale consistency test rather than a population-level fit.
The localized branch gives
$\begin{gathered}{\rm median}\left(\frac{R_{\rm gap}}{R_{\rm ring}}\right)=0.818,\\ 16{\rm--}84\%:\;0.677{\rm--}0.906.\end{gathered}$ (31)
The representative observed ratios for Hoag’s Object, UGC 4599, and PGC 1000714 are approximately
$0.890,\qquad 0.780,\qquad 0.717,$ (32)
respectively (Schweizer et al., 1987; Finkelman et al., 2011; Finkelman & Brosch, 2011; Pakdil et al., 2017). All three fall within the central model interval. JO171, whose ring has been interpreted as externally formed and is now being reshaped by ram-pressure stripping in a cluster environment, has
$\left(R_{\rm gap}/R_{\rm ring}\right)_{\rm JO171}\simeq 0.325$ (33)
using the ring extent and inner rise radius from the MUSE analysis (Moretti et al., 2018); it lies outside the clean shell-localized interval. This is not a decisive exclusion test for JO171’s formation history, but it is consistent with treating JO171 as an environmentally processed comparison object rather than a member of the isolated shell-supported branch.
With $R_{\rm ring}^{\rm obs}=17\,{\rm kpc}$ as a fiducial Hoag-like calibration, the localized branch gives median physical scales
$\begin{gathered}R_{\rm gap}\simeq 13.9\,{\rm kpc},\qquad R_{\rm shell}\simeq 13.9\,{\rm kpc},\\ \Delta_{\rm phys}\simeq 0.94\,{\rm kpc}.\end{gathered}$ (34)
The near equality of $R_{\rm gap}$ and $R_{\rm shell}$ is the physical expression of $x_{\rm g}\simeq x_{0}$: the depleted gap should trace the settled shell radius.
[…]
Alt text: Probability-density plot of the model gap-to-ring ratio with observed markers for Hoag’s Object, UGC 4599, PGC 1000714, and JO171; the first three overlap the model interval and JO171 lies below it.
Formation interpretation and outlook.
The advantage of the present construction is not that it replaces hydrodynamical or cosmological formation modeling. Its role is to isolate the radial selection step required once an old compact core and a later settled gas reservoir are present. A plausible history is the following. A main halo forms in a relatively low-density environment while remaining connected to a cold-gas supply channel (Sancisi et al., 2008; Kereš et al., 2005; Dekel et al., 2009). Early low-angular-momentum gas falls deeper into the halo, cools efficiently, and forms the old central spheroidal component. Mildly anisotropic early accretion, gentle relaxation, or weak mergers may leave this core slightly triaxial without producing a strongly barred disk.
Later, higher-angular-momentum gas from the same large-scale supply can enter the halo directly or through gas-rich, star-poor subhalos. Such subhalos may remain optically dark if their potential wells are shallow, their gas surface densities are low, or the ultraviolet background suppresses star formation (Efstathiou, 1992). If they enter the main halo on large impact parameter orbits, their H i can be stripped without leaving an obvious stellar tidal tail. The stripped gas inherits the orbital angular momentum of the infall, dissipates radial energy, and partially circularizes into an extended finite-radius reservoir. Continued supply from the same channel can maintain an H i structure larger than the optical ring and can produce a mild warp if the later angular- momentum distribution is broader than that of the initial ring-forming event.
In this interpretation, the Gaussian shell in Eq. (4) is the lowest-order radial imprint of the settled H i reservoir. It is not a literal self-consistent gravitational potential of the gas. A weak quadrupole from a mildly triaxial central body may modulate the reservoir density and help selected annuli cross a star-formation threshold, while remaining too weak to destroy the near-circular outer ring. The boundary-supported max–min branch supplies the missing selection step: it separates the outer settling basin from the compact inner component and keeps the intermediate region underpopulated. This is why the absence of an interior $x_{\rm core}$ is not a defect of the model; the core is a pre-existing boundary state, while the shell organizes only the newly accreted material outside it.
The mechanism is therefore complementary to both collisional-wave and bar-resonance pictures. It does not require a recent intruder launching an expanding density wave, nor a strong long-lived bar. It also differs from a fully self-gravitating massive-ring orbital model: the shell is an effective localized control deformation associated with settled accreted material. The immediate tests are correspondingly direct. Clean Hoag-type systems should show an old central population, a younger or actively star-forming outer ring, a gap radius correlated with the H i shell or reservoir radius, and a ring with small radial motion relative to its azimuthal motion. Environmentally processed systems, such as JO171, need not lie on the same clean shell-localized scaling relation.
Several extensions are direct. A multi-shell control potential,
$\Phi(x)=-\frac{\mu}{x}+\sum_{i}\epsilon_{i}\exp\!\left[-\frac{(x-x_{0,i})^{2}}{2\Delta_{i}^{2}}\right],$ (35)
would test whether double-ring systems such as PGC 1000714 can be understood as sequential accretion-supported shells. A dynamical follow-up should integrate ensembles of dissipative gas parcels in the time-dependent shell potential, include angular-momentum dispersion and shell growth, and measure whether
$\frac{\Sigma(x_{\rm r})}{\Sigma(x_{\rm g})}>C_{\rm gap},\qquad\frac{\Delta x_{\rm ring}}{x_{\rm r}}<q_{\rm ring},\qquad\frac{\langle|v_{R}|\rangle_{\rm ring}}{\langle v_{\varphi}\rangle_{\rm ring}}<\eta.$ (36)
These diagnostics would distinguish a dissipatively settled ring from an expanding collisional wave and would connect the boundary-supported control mechanism to observable H i, stellar-age, and kinematic maps.
Acknowledgements
We thank the authors of the observational studies cited here for making Hoag-like systems a quantitatively testable morphology class. We also thank the developers of the open-source Python and tools used in this work.
Data Availability
The observational quantities used in this work are taken from the cited literature. The Monte Carlo catalogues and plotting scripts underlying this article will be shared by the corresponding authors upon reasonable request.